# Weight Ensembling Improves Reasoning in Language Models

### v1

@@ Abstract
We investigate a failure mode that arises during the training of reasoning models, where the diversity of generations begins to collapse, leading to suboptimal test-time scaling. Notably, the Pass@1 rate reliably improves during supervised finetuning (SFT), but Pass@k rapidly deteriorates. Surprisingly, a simple intervention of interpolating the weights of the latest SFT checkpoint with an early checkpoint, otherwise known as WiSE-FT, almost completely recovers Pass@k while also improving Pass@1. The WiSE-FT variant achieves better test-time scaling (Best@k, majority vote) and achieves superior results with less data when tuned further by reinforcement learning. Finally, we find that WiSE-FT provides complementary performance gains that cannot be achieved only through diversity-inducing decoding strategies, like temperature scaling. We formalize a bias-variance tradeoff of Pass@k with respect to the expectation and variance of Pass@1 over the test distribution. We find that WiSE-FT can reduce bias and variance simultaneously, while temperature scaling inherently trades-off between bias and variance.
@@ 1 Introduction
Unfortunately, while the standard training pipeline of supervised finetuning (SFT) followed by reinforcement learning (RL) dependably improves $\mathrm{Pass@1}$ for reasoning, $\mathrm{Pass@K}$ tends to drop early into finetuning (Cobbe et al., 2021; Chow et al., 2024a; Chen et al., 2025). This mismatch arises from a symptom of finetuning called diversity collapse, where overtuned models yield less diverse generations. This is detrimental to $\mathrm{Pass@K}$ since the model is wasting $K$ attempts on only a handful of guesses. In fact, by analyzing the model’s error rate ($1-\mathrm{Pass@1}$) across the test distribution, we derive a $\mathbf{\mathrm{Pass@K}}$ bias-variance trade-off. To improve expected test $\mathrm{Pass@K}$, one can either reduce the bias which is the expected error rate or how much the model’s error rate varies across problems. The latter term is connected to diversity — more diversity allows models to hedge and do uniformly well across all test questions. In particular, during SFT, $\mathrm{Pass@1}$ improves (bias $\downarrow$) at the cost of diversity collapse (variance $\uparrow$).
Surprisingly, common ways of alleviating diversity collapse, such as early stopping SFT at peak $\mathrm{Pass@K}$ or decoding with high temperature, suffer from the reverse trade-off: diversity improves (variance $\downarrow$) at the cost of overall $\mathrm{Pass@1}$ degrading (bias $\uparrow$). Thus, in this paper we are concerned with a central question:
@@
In summary, we provide a comprehensive analysis of how reasoning models suffer from diversity collapse during SFT and its negative downstream impact during RL and test-time scaling. We first discuss our WiSE-FT findings in §4. Motivated by this discovery, we investigate two fundamental questions. First, we conduct a theoretical investigation of diversity collapse during SFT and RL of reasoning models in §5. Diversity collapse not only impacts the model’s ability to attempt different guesses. In fact, we make an even stronger observation — the generations of reasoning models converge towards a single reasoning trace for each test question. We prove that standard RL algorithms (i.e., REINFORCE and GRPO) fail to recover lost diversity in a simplified discrete bandit setting.
@@ 2 Related Works
The standard pipeline for enhancing reasoning in LLMs involves an initial phase of supervised fine-tuning (SFT) followed by reinforcement learning (RL) (Guo et al., 2025; Setlur et al., 2024). SFT is critical for instilling interpretable and readable reasoning chains and ensuring that the model adheres to a consistent rollout templates  (Guo et al., 2025). However, a number of recent works have identified critical pitfalls of SFT that hinders the model’s ability to explore and ultimately it’s overall problem solving ability. Notably, Cobbe et al. (2021) observe diversity collapse when finetuning on GSM8k training dataset, during which the Pass@1 continuously improves whereas Pass@k starts to fall shortly into the training. This is not desirable as diverse sampling at inference is important for test-time scaling using majority voting (Wang et al., 2023) or reward model guided search (Setlur et al., 2024; Beeching et al., 2024). Yeo et al. (2025); Chu et al. (2025) attribute this behavior to overfitting, memorization of samples and overfixation to a template style leading to reduced generalization. In our work, we corroborate similar findings and propose ensembling over the course of SFT as a strong mitigation to the diversity collapse and analyze the same.
@@
Given the importance of diversity for effectively scaling inference-time compute, several recent works have proposed auxiliary finetuning objectives and decoding strategies to mitigate diversity collapse. Li et al. (2025) regularize the SFT process using a game-theoretic framework that encourages sparse updates, thereby preserving output diversity.  Zhang et al. (2024b) directly optimizes for diversity during finetuning. Other approaches modify the finetuning procedure to directly optimize for Best-of-N sampling at inference time (Chow et al., 2024b; Sessa et al., 2024; Chen et al., 2025). Another line of work focuses on inference-time decoding, explicitly encouraging diverse solutions through modified beam search strategies (Vijayakumar et al., 2018; Olausson et al., 2024; Chen et al., 2024; Beeching et al., 2024). Li et al. (2023) improve diversity during parallel decoding by appending curated prompts to the input. In formal reasoning settings (e.g., Lean), methods such as Monte Carlo tree search have been used to diversify intermediate reasoning steps, as demonstrated in AlphaProof (AlphaProof and AlphaGeometry teams, 2024). In this work, we identify a simple and complementary intervention during the finetuning process to maintain the diversity of generations.
@@ 3.1 Pass@k, Best@k, and Majority Vote
where $\rho_{x}=P(\hat{y}=y\mid x,f,D)$ is the accuracy or marginal probability of sampling the ground truth answer. Then $(1-\rho_{x})^{K}$ is the probability that all $K$ guesses are incorrect. Note that $\rho_{x}$ is also equal to $\mathrm{Pass@1}$. We will refer to $\mathrm{Pass@1}$ as $\rho_{x}$ interchangeably in our paper.
@@
Notably, $\mathrm{Pass@K}$ is equivalent to $\mathrm{Best@K}$ using a perfect ORM verifier. As we will observe, WiSE-FT achieves higher $\mathrm{Pass@1}$ $\mathrm{Pass@K}$ and this directly translates to achieving better $\mathrm{Best@K}$ both with an ORM verifier and by Majority Vote.
@@ 3.3 Training and Evaluation Pipeline
The majority of our experiments are conducted on Gemma-2-2B and Qwen-2.5-0.5B . We perform SFT on a 30K subset of rephrased augmentations of GSM8k (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) in MetaMath40k (Yu et al., 2023) for 1710 steps or 10 epochs. We then continue finetuning on another 30K subset of rephrased training questions from MetaMath using Group Relative Policy Optimization (GRPO) with a binary reward of the correctness of the model’s final answer. Finally, we evaluate models on GSM8K and MATH500, respectively. To estimate the true $\mathrm{Pass@K}$ and $\mathrm{Pass@1}$ marginalized over the distribution of sampled traces, we sample 100 reasoning traces per test example and average over them to estimate $\mathrm{Pass@1}$, i.e. $\rho_{x}$. Then to calculate $\mathrm{Pass@K}$, we use the theoretical formula $1-(1-\rho_{x})^{K}$ in Equation 1. Unless noted otherwise, we employ a naive decoding strategy with top-p threshold $0.9$, temperature $T=0.8$, and top-k $K=50$.
@@ 4 Improving Diverse Reasoning Capabilities by WiSE-FT
We first carefully track $\mathrm{Pass@K}$ for $K\in\{1,4,32\}$ across the SFT trajectory of Qwen-2.5-0.5B and Gemma-2-2B . Similar to findings from Cobbe et al. (2021); Chen et al. (2025), we observe that $\mathrm{Pass@1}$ continues to improve with longer SFT, whereas for larger $K=4,32$, $\mathrm{Pass@K}$ tends to peak much earlier on in training (in Figure 1, 16, and 18). In other words, while later SFT checkpoints achieve higher $\mathrm{Pass@1}$, earlier SFT checkpoint achieve higher $\mathrm{Pass@K}$. This tradeoff in model selection is not ideal downstream for test-time scaling.
@@ 4.1 General Purpose Reasoning Models
So far we have studied the effect of WiSE-FT on models tuned on reasoning data for the same specific reasoning task (e.g., train on GSM8k and evaluate on GSM8k). We’ve additionally tested how well our findings generalize to models trained on general purpose reasoning datasets and tested on a out-of-distribution reasoning task. We take Qwen2.5-7B-Instruct and SFT for 5 epochs on OpenThoughts-114k, a high-quality synthetic dataset of math, science, and coding questions paired with DeepSeek-R1 completions, then evaluate its performance on AIME24 competition problems (with ASY code for figures from Muennighoff et al. (2025)). In this setting, the $\mathrm{Pass@K}$ trends during SFT on is more subtle. We still observe diversity collapse in Figure 12, but the affect is not strong enough for $\mathrm{Pass@K}$ to drop back down. However, we observe that the rate at which Pass@K improves for $K\in\{16,32\}$ slows down early while Pass@1 grows at a constant rate (Figure 10). We then perform WiSE-FT between the final and earlier checkpoint with higher diversity. We choose early checkpoint at epoch 3 where improvements in $\mathrm{Pass@K}$ begin to slow.
@@ 5 Diversity Collapse during Finetuning
In previous sections we alluded to the phenomenon where $\mathrm{Pass@K}$ decreases because SFT induces diversity collapse in reasoning traces. To verify this hypothesis, we sample 100 traces per test GSM8k problem and measure diversity using three metrics:
@@ 5.1 Theoretical Discussion of Diversity Collapse During SFT and RL
We further prove why applying reinforcement learning to a low-diversity policy yields suboptimal results—and sometimes even further diversity collapse—in a discrete bandit setting (see Figure 5). In this scenario, we assume there exist $K$ equally good arms, corresponding to a set of successful strategies, and one bad arm that the policy should learn to avoid. We show two key results in this setting:
@@
Figure 6: Histogram of error $\mathbf{1-\rho_{x}}$ of Gemma-2-2B SFT checkpoints across GSM8k test. SFT progressively decreases bias but increases variance of $1-\mathrm{Pass@1}$ across the test distribution due to diversity collapse, causing $\mathrm{Pass@K}$ to fall. Applying Wise-FT reduces both bias and variance, but temperature scaling trades off decreasing variance with increased bias.
@@ 6.1 Diversity Collapse leads to Bimodal Pass@1 Distribution
In Figure 6(b), we plot the distribution of $1-\rho_{x}$, estimated using 100 sampled traces, over GSM8K test. We notice two trends with longer SFT. First, biases decreases, i.e., the expectation of this distribution shifts towards 0. However, the distribution becomes increasingly bimodal with the densities converging towards the two extremes 0 and 1. As a result, the variance increases with longer SFT. This increase in variance directly explains the drop in Pass@k.
The bimodality of the $1-\rho_{x}$ distribution means that the $\mathrm{Pass@1}$ of any test problem is either very high or very low. Interestingly, one explanation for the increased bimodality of the distribution of $1-\rho_{x}$ is in fact when models suffer from diversity collapse causing a particular guess to be oversampled. In other words, for a test problem, if the model places high probability on an incorrect guess, $\mathrm{Pass@1}$ is very low. On the other hand, if the model places high probability on the correct guess, $\mathrm{Pass@1}$ is very high. We illustrate this relationship in Figure 6(a). All in all, $\mathrm{Pass@K}$ can be improved in two ways — either reduce bias by improving $\mathrm{Pass@1}$ or reduce variance by increasing diversity.
@@
While we’ve proposed inducing diversity by WiSE-FT, another common alternative for inducing diversity is temperature scaling the logits. High temperature smoothens the logits allowing the model to more likely sample low probability tokens. In Figure 1, we see that while high temperatures indeed improve $\mathrm{Pass@K}$, the $\mathrm{Pass@K}$ at any SFT timestep notably never reaches the $\mathrm{Pass@K}$ of our final WiSE-FT model. If temperature scaling also increases diversity, why does WiSE-FT strictly outperform sampling with high temperature? In Figure 6(b), we plot the distribution of $1-\rho_{x}$ if we sample from the last SFT checkpoint with high temperature $T=1.6$. As expected, we see that the model reasons more diversely. This smoothens the bimodal peaks and reduces the variance. However, the average accuracy of the model generations also degrades, causing the bias goes back up. We suspect bias-variance tradeoff is inherent in diversity-inducing decoding approaches. For example, min-p (Nguyen et al., 2024) combines temperature scaling with adaptive thresholding to not sample outlier tokens. However, this additional control is unable to reduce bias. Surprisingly, WiSE-FT uniquely manages to reduce both bias and variance.

### v3

@@ Abstract
We investigate a failure mode that arises during the training of reasoning models, where the diversity of generations begins to collapse, leading to suboptimal test-time scaling. Notably, the Pass@1 rate reliably improves during supervised finetuning (SFT), but Pass@k rapidly deteriorates. Surprisingly, a simple intervention of interpolating the weights of the latest SFT checkpoint with an early checkpoint, otherwise known as WiSE-FT, almost completely recovers Pass@k while also improving Pass@1. The WiSE-FT variant achieves better test-time scaling (Best@k, majority vote) and achieves superior results with less data when tuned further by reinforcement learning. Finally, we find that WiSE-FT provides complementary performance gains that cannot be achieved only through diversity-inducing decoding strategies, like temperature scaling. We formalize a bias-variance tradeoff of Pass@k with respect to the expectation and variance of Pass@1 over the test distribution. We find that WiSE-FT can reduce bias and variance simultaneously, while temperature scaling inherently trades off between bias and variance.
@@ 1 Introduction
Unfortunately, while the standard training pipeline of supervised finetuning (SFT) followed by reinforcement learning (RL) dependably improves $\mathrm{Pass@1}$ for reasoning, $\mathrm{Pass@K}$ tends to drop early into finetuning (Cobbe et al., 2021; Chow et al., 2024a; Chen et al., 2025). This mismatch arises from a symptom of finetuning called diversity collapse, where overtuned models yield less diverse generations. This is detrimental to $\mathrm{Pass@K}$ since the model wastes $K$ attempts on only a handful of guesses. In fact, by analyzing the model’s error rate i.e., $1-\mathrm{Pass@1}$, across the test distribution, we derive a $\mathbf{\mathrm{Pass@K}}$ bias-variance trade-off. To improve expected test $\mathrm{Pass@K}$, one can either reduce the bias which is the expected error rate or how much the model’s error rate varies across problems. The latter term is connected to diversity — more diversity allows models to hedge and do uniformly well across all test questions. In particular, during SFT, $\mathrm{Pass@1}$ improves (bias $\downarrow$) at the cost of diversity collapse (variance $\uparrow$).
Surprisingly, common ways of alleviating diversity collapse, such as early stopping at peak $\mathrm{Pass@K}$ or decoding with high temperature, suffer from the reverse trade-off: diversity improves (variance $\downarrow$) at the cost of overall $\mathrm{Pass@1}$ degrading (bias $\uparrow$). Consequently, in this paper we are concerned with a central question:
@@
In summary, we provide a comprehensive analysis of how reasoning models suffer from diversity collapse during SFT and its negative downstream impact during RL and test-time scaling. We first discuss our WiSE-FT findings in §4. Motivated by this discovery, we investigate two fundamental questions. First, we investigate diversity collapse during SFT and RL of reasoning models in §5. Diversity collapse not only impacts the model’s ability to attempt different guesses. In fact, we make an even stronger observation — the generations of reasoning models converge towards a single reasoning trace for each test question. We theoretically prove that standard RL algorithms (i.e., REINFORCE and GRPO) fail to recover lost diversity in a simplified discrete bandit setting.
@@ 2 Related Works
The standard pipeline for enhancing reasoning in LLMs involves an initial phase of supervised fine-tuning (SFT) followed by reinforcement learning (RL) (Guo et al., 2025; Setlur et al., 2024). SFT is critical for instilling interpretable and readable reasoning chains and ensuring that the model adheres to a consistent rollout templates  (Guo et al., 2025). However, a number of recent works have identified critical pitfalls of SFT that hinders the model’s ability to explore and ultimately it’s overall problem solving ability. Notably, Cobbe et al. (2021) observe diversity collapse when finetuning on GSM8k training dataset, during which the Pass@1 continuously improves whereas Pass@k starts to fall shortly into the training. Similar diversity collapse phenomenon also exists in the self-improvement setting with SFT (Song et al., 2024), and is theoretically investigated as the sharpening effect (Huang et al., 2024). This is not desirable as diverse sampling at inference is important for test-time scaling using majority voting (Wang et al., 2023) or reward model guided search (Setlur et al., 2024; Beeching et al., 2024). Yeo et al. (2025); Chu et al. (2025) attribute this behavior to overfitting, memorization of samples and overfixation to a template style leading to reduced generalization. In our work, we corroborate similar findings and propose ensembling over the course of SFT as a mitigation strategy.
@@
Given the importance of diversity for effectively scaling inference-time compute, several recent works have proposed auxiliary finetuning objectives and decoding strategies to mitigate diversity collapse. Li et al. (2025) regularize the SFT process using a game-theoretic framework that encourages sparse updates, thereby preserving output diversity.  Zhang et al. (2024b) directly optimizes for diversity during finetuning. Other approaches modify the finetuning procedure to directly optimize for Best-of-N sampling at inference time (Chow et al., 2024b; Sessa et al., 2024; Chen et al., 2025). Another line of work focuses on inference-time decoding, explicitly encouraging diverse solutions through modified beam search strategies (Vijayakumar et al., 2018; Olausson et al., 2024; Chen et al., 2024; Beeching et al., 2024). Li et al. (2023) improve diversity during parallel decoding by appending curated prompts to the input. In formal reasoning settings e.g., Lean, methods such as Monte Carlo tree search have been used to diversify intermediate reasoning steps, as demonstrated in AlphaProof (AlphaProof and AlphaGeometry teams, 2024). In this work, we identify a simple and complementary intervention during the finetuning process to maintain the diversity of generations. We especially care about enforcing diversity while preserving the overall accuracy of generations.
@@ 3.1 Pass@k, Best@k, and Majority Vote
where $\rho_{x}=P(\hat{y}=y\mid x,f,D)$ is the Pass@1 or marginal probability of sampling the ground truth answer. Then $(1-\rho_{x})^{K}$ is the probability that all $K$ guesses are incorrect. We will refer to $\mathrm{Pass@1}$ as $\rho_{x}$ interchangeably in our paper.
@@
Notably, $\mathrm{Pass@K}$ is equivalent to $\mathrm{Best@K}$ using a perfect ORM verifier. As we will observe, WiSE-FT achieves both higher $\mathrm{Pass@1}$ and $\mathrm{Pass@K}$ and this directly translates to achieving better $\mathrm{Best@K}$ with an ORM verifier and by Majority Vote.
@@ 3.3 Training and Evaluation Pipeline
The majority of our experiments are conducted on Gemma-2-2B and Qwen-2.5-0.5B . We perform SFT on a 30K subset of rephrased augmentations of GSM8k (Cobbe et al., 2021) and MATH (Hendrycks et al., 2021) in MetaMath40k (Yu et al., 2023) for 1710 steps or 10 epochs. We then continue finetuning on another 30K subset of rephrased training questions from MetaMath using Group Relative Policy Optimization (GRPO) with a binary reward of the correctness of the model’s final answer. Finally, we evaluate models on GSM8K and MATH500, respectively. To estimate the true $\mathrm{Pass@K}$ and $\mathrm{Pass@1}$ marginalized over the distribution of sampled traces, we sample 100 reasoning traces per test example and average over them to estimate $\mathrm{Pass@1}$, i.e. $\rho_{x}$. Then to calculate $\mathrm{Pass@K}$, we use the theoretical formula $1-(1-\rho_{x})^{K}$ in Equation 1. Unless noted otherwise, we employ a naive decoding strategy with top-p threshold $0.9$, temperature $T=0.8$, and top-k with $K=50$.
@@ 4 Improving Diverse Reasoning Capabilities by WiSE-FT
We first carefully track $\mathrm{Pass@K}$ for $K\in\{1,4,32\}$ across the SFT trajectory of Qwen-2.5-0.5B and Gemma-2-2B . Similar to findings from Cobbe et al. (2021); Chen et al. (2025), we observe that $\mathrm{Pass@1}$ continues to improve with longer SFT, whereas for larger $K=4,32$, $\mathrm{Pass@K}$ tends to peak much earlier on in training (in Figure 1, 17, and 19). In other words, while later SFT checkpoints achieve higher $\mathrm{Pass@1}$, earlier SFT checkpoint achieve higher $\mathrm{Pass@K}$. This tradeoff in model selection is not ideal downstream for test-time scaling.
@@ 4.1 General Purpose Reasoning Models
So far we have studied the effect of WiSE-FT on models tuned on reasoning data for the same specific reasoning task (e.g., train on GSM8k and evaluate on GSM8k). We’ve additionally tested how well our findings generalize to models trained on general purpose reasoning datasets and tested on a out-of-distribution reasoning task. We take Qwen2.5-7B-Instruct and SFT for 5 epochs on OpenThoughts-114k, a high-quality synthetic dataset of math, science, and coding questions paired with DeepSeek-R1 completions, then evaluate its performance on AIME24 competition problems (with ASY code for figures from Muennighoff et al. (2025)). In this setting, the $\mathrm{Pass@K}$ trends during SFT on is more subtle. We still observe diversity collapse in Figure 12, but the affect is not strong enough for $\mathrm{Pass@K}$ to drop back down. However, we observe that the rate at which Pass@K improves for $K\in\{16,32\}$ slows down early while Pass@1 grows at a constant rate (Figure 10). We then perform WiSE-FT between the final and earlier checkpoint with higher diversity. We choose early checkpoint at epoch 3 where improvements in $\mathrm{Pass@K}$ begin to slow. Similarly, we observe that WiSE-FT improves both Pass@1 and Pass@K in Figure 2.
@@ 5 Diversity Collapse during Finetuning
In previous sections we alluded to the phenomenon where $\mathrm{Pass@K}$ decreases because SFT and RL induces diversity collapse in reasoning traces. To verify this hypothesis, we sample 100 traces per test GSM8k problem and measure diversity using three metrics:
@@ 5.1 Theoretical Discussion of Diversity Collapse During SFT and RL
We further prove why applying reinforcement learning to a low-diversity policy yields suboptimal results—and sometimes even exacerbates diversity collapse—in a discrete bandit setting (see Figure 5). In this scenario, we assume there exist $K$ equally good arms, corresponding to a set of successful strategies, and one bad arm that the policy should learn to avoid. We show two key results in this setting:
@@
Figure 6: Histogram of error $\mathbf{1-\rho_{x}}$ of Gemma-2-2B SFT checkpoints across GSM8k test. SFT progressively decreases bias but increases variance of error i.e., $1-\mathrm{Pass@1}$, across the test distribution, causing $\mathrm{Pass@K}$ to fall. Applying Wise-FT reduces both bias and variance, but temperature scaling trades off decreasing variance with increased bias.
@@ 6.1 Diversity Collapse leads to Bimodal Pass@1 Distribution
In Figure 6(b), we plot the distribution of error $1-\rho_{x}$, estimated using 100 sampled traces, over GSM8K test examples. We notice two trends with longer SFT. First, bias decreases, i.e., the expected error shifts towards 0. However, the distribution becomes increasingly bimodal with the densities converging towards the two extremes 0 and 1. As a result, the variance increases with longer SFT. This increase in variance directly explains the drop in Pass@k.
The bimodality of the $1-\rho_{x}$ distribution means that the $\mathrm{Pass@1}$ of any test problem is either very high or very low. Interestingly, one explanation for the increased bimodality of the distribution of $1-\rho_{x}$ is in fact when models suffer from diversity collapse. In other words, a particular guess to be oversampled for each test problem. If the model places high probability on an incorrect guess, $\mathrm{Pass@1}$ is very low. On the other hand, if the model places high probability on the correct guess, $\mathrm{Pass@1}$ is very high. We illustrate this relationship in Figure 6(a). All in all, $\mathrm{Pass@K}$ can be improved in two ways — either reduce bias by improving $\mathrm{Pass@1}$ or reduce variance by increasing diversity.
@@
While we’ve proposed inducing diversity by WiSE-FT, another common alternative for inducing diversity is temperature scaling the logits. High temperature smoothens the logits allowing the model to more likely sample low probability tokens. In Figure 1, we see that while high temperatures indeed improve $\mathrm{Pass@K}$, the $\mathrm{Pass@K}$ at any SFT timestep notably never reaches the $\mathrm{Pass@K}$ of our final WiSE-FT model. If temperature scaling also increases diversity, why does WiSE-FT strictly outperform sampling with high temperature? In Figure 6(b), we plot the distribution of $1-\rho_{x}$ if we sample from the last SFT checkpoint with high temperature $T=1.5$. As expected, we see that the model reasons more diversely. This smoothens the bimodal peaks and reduces the variance. However, the average accuracy of the model generations also degrades, causing the bias goes back up. We suspect bias-variance tradeoff is inherent in diversity-inducing decoding approaches. For example, min-p (Nguyen et al., 2024) combines temperature scaling with adaptive thresholding to not sample outlier tokens. However, this additional control is unable to reduce bias (Figure 16). Surprisingly, WiSE-FT uniquely manages to reduce both bias and variance.